An experiment in a general chemistry laboratory calls for a 2.00-Msolution of HCl. How many mL of 11.9M HCl would be required to make 250 mL of 2.00 MHCl?

Please show me the answer step by step

Thanks..
Answer #1

Number of moles = 2.00 M * 0.25 mL = 0.5 mole

Molarity = number of moles / volume of solution in L

11.9 M = 0.5 mole / volume of solution in L

volume of solution in L = 0.5/11.9 = 0.042 L

1L = 1000 mL

0.042 L = 42.01 mL

Therefore, 42.01 mL of 11.9M HCl solution is required.

500 = 11.9 M

V2 = 42.01 mL